A compressed air tank contains 5 kg of air at a temperature of 80 °C. A gauge pressure meter on the tank reads 300 kPa. Determine the volume of the tank and state the assumption made


1
Expert's answer
2021-07-17T04:52:36-0400

Consider Ideal gas law, express "\rho" as mass per volume

"p = \rho TR"


"p = \frac{m}{V}TR"


"V = \frac{m}{p}TR = \frac{ 5 kg.353K.287.05\frac{j}{kgK}}{(300+101.6)10^3 pa}= 1.2625 m^3"